Assume that the weekly demand for Ford car sales follows a normal distribution with mean 50,000 cars and standard deviation 14,000 cars.
a) There is a 1% chance that Ford will sell more than what number of cars during the next year?
b) What is the probability that Ford will sell between 2.4 and 2.7 million cars during the next year?

Answer by stanbon(75887)

Assume that the weekly demand for Ford car sales follows a normal distribution with mean 50,000 cars and standard deviation 14,000 cars.
----------
a) There is a 1% chance that Ford will sell more than what number of cars during the next year?
Find the z-value with a left tail of 99%: z = 2.3263
----
Find the corresponding x(car sales number) using x = z*s+u
---
# of car sales = 2.3263*14000+50000 = 82569
================================================

b) What is the probability that Ford will sell between 2.4 and 2.7 million cars during the next year?
Find the z-values for 2.4 mil and 2.7 mil ; then find the probability
that lies between those z-values.

========================================
Cheers,
Stan H.